Context
At the SOWG Kick-off Meeting on several changes to the plan were already mentioned to the team. We record them here, along with the progressive optimisations of the plan for coordination and TM usage.
Clarifications to SOOPs
- 2022-10-16 to 27: R_SMALL_MRES_MCAD_AR-Long-Term + R_SMALL_HRES_MCAD_AR-Heating: It was clarified that both SOOPs can run on the same target AR, to simplify the plan and the decisions to be made.
Proposals for updates:
Those with a symbol indicate a request for TM beyond the MLP science plan.
- 2022-10-21 to 25: R_SMALL_HRES_MCAD_AR-Heating should consist of 5 instances (not 3) instances of the same SOOP, 5-hours runs each, led by SPICE.
changes made in SOOP Kitchen: the placeholder SOOP labels L_TEMPORARY were removed and new 5-hour long SOOPs of type AR-Heating created based on the SPICE observations (placeholders still, to be updated during SOWG)
resulting version: v26
- 2022-10-29 to 31: R_SMALL_HRES_MCAD_AR-Heating : SPICE clarified these are 3 short versions of the new SOOP, only 2 hours long each.
Labels still to be updated with correct SOOP name.Done. EUI observations also included in the SOOP instances.
2022-10-12 to 13: R_FULL_HRES_HCAD_Density-Fluctuations : Metis proposes to increase both the spatial and temporal resolution of Metis observations during this SOOP.
This would account to an increase in TM of 2900 MiB. It would cause a high peak in Metis data generation in the period Oct 12 19:00 to Oct 13 12:00, i.e. 3700 MiB over 27 hrs.
The extra TM could be partially compensated by gaps for WOLs during the Coronal Dynamics SOOP earlier.
SOC to analyse feasibility of this increase. https://solarorbiter.esac.esa.int/soopkitchen/#/planning/child-plan/LTP09_Metis_scenarios baseline 1Will report at SOWG.
2022-10-12 to 16: EUI brings up that in this period, 4 burst campaigns with HRI are planned. The two first ones, associated with R_SMALL_HRES_HCAD_RS-burst have representative TM in the plan, but the last 2 campaigns for the SPICE raster (CC_OFFPOI_OOF) and R_SMALL_HRES_MCAD_Polar-Observations are underrepresented in terms of EUI TM.
There are 2 options, to be explored:- run all 4 campaigns at 5s cadence for HRI/EUV → significant TM increase reflected in v10 of EUI child plan
Ideally:- enhance the HRIEUV support for the west-limb off-pointing to 1440 MiB (+720 MiB) making it 2 hours at 5s cadence (also reflected in v10 of EUI child plan)
- for the polar campaign, increase to HRI's maximum cadence at 2s. In 1 hour that would give ideally 1800 MiB (+1440 MiB more than currently) (reflected in v12 of EUI child plan)
- skip 1 or more of these campaigns, potentially the 12 Oct burst sequence at disk centre
- run all 4 campaigns at 5s cadence for HRI/EUV → significant TM increase reflected in v10 of EUI child plan
- 2022-10-16 to 27: R_SMALL_MRES_MCAD_AR-Long-Term: EUI confirms it can support the SOOP with regular (daily) bursts of HRI observations, however it would be preferred to run HRI-EUV at 5s cadence instead of 10s in order to gain data quality and reduce degradation of the instrument.
Implemented independently of Option 4 above up to 2022-10-27T1900, in v 40 of the EUI child plan.
- If no more TM would be available, these are the options:
- reduce the 30mins burst campaigns to 15 mins
- keep 30mins slots but reduce the frequency (preferred by SOOP coordinator S. Parenti
SOC to analyse whether any more TM would be available in this time period
SOOP coordinators & SOWG to discuss preferred way forward
- If no more TM would be available, these are the options:
- 2022-11-06: COORD_CALIBRATION: SOC informed the SOWG during kick-off that the "Full Disk Mosaic for connection science" (aka Crazy Mosaic) overlaps with a pass during which we cannot update the pointing.
The SOOP coordinator confirmed that the SOOP can be shifted earlier, to avoid the PASS and if possible also the preceding HGA_PT.
POINT_PATTERN event has been placed earlier starting at 2022-11-06 9:40. (Observations still to be shifted accordingly, in coordination with AR-long-term SOOP coordinator) - LTP9 up to 2022-10-26: PHI asks whether it would be possible to reshuffle the PHI flushes in the plan in order to allow formatting a partition of the PHI internal memory on 2022-10-26. In order to empty that partition, an extra volume of 2300MiB would need to be flushed before October 26 (shifting later flushes earlier, and spreading them out in, e.g., daily flushes). The total data volume over the whole LTP would not increase.
SOC to investigate whether this scenario is possible:
- see child plan https://solarorbiter.esac.esa.int/soopkitchen/#/planning/child-plan/LTP09_PHI_scenarios (bookmarked v7)
- extra 2300MiB spread over first month of LTP9 = 76.6 MiB/day → daily flushes in addition to the existing ones
- this volume got subtracted from the 4 big flushes on 29, 30, 31 Oct and 1 Nov (-575 each)
- simulation ran with updated itls (v396 of parent + phi updated flushes) and original EFECS
- resulting ssmm fill states compared with LTP9 plan baseline 1 (v93) -
SOC will report at SOWG
RPW would like to require 4 TAC windows during LTP09, each of 90 MiB, and for a duration of 24h. The dates are informative and can be moved by one day or two.
There will allow us to run BIAS tests and measurements at different distances from the Sun to study the impact on the 50 KHz interference.Dates proposal :
- at the very beginning of LTP09, asap according to the superior conjunction constraints
- ~ Oct 25
- ~ Nov 15
- ~ Dec 5
SOC to investigate this scenario.Placed in main plan: cannot add events in Child Plan. Some timings adjusted because of when PASS, MAINT and con events fall, and when STPs start and end. RPW informed.
Proposal from SPICE agreed by affected SOOP Coordinators (Susanna, Gherardo) :
Connection Mosaic (on 13-16 October 2022) in the past few days. This would affect also the West Limb Quadrature (CC_OFFPOI_OOF), Polar Observations and the beginning of AR Long Term SOOPs. Reduces the total length of the AR Long Term SOOP by 11 hours, keeps a reasonable pattern for the Connection Mosaic given the WOL and PASS constraints.
L_SMALL_MRES_MCAD_Connection-Mosaic would run from 15:00 on 13 October to 00:00 on 16 October. The proposed pointings duration would be as follows,
Pointing 1: 13/10/2022 16:00implemented by SOC with small changes highlighted in blue:
15:00-00:00 (8hr)
Pointing 2: 14/10/2022 03:10-13:15 (10hr 5min)
Pointing 3: 14/10/2022 13:20-23:00 (9hr 40min)
Pointing 4: 14/10/2022 23:05-06:05 (7hr)
Pointing 5: 15/10/2022 06:10-13:15 (7hr 5min)
Pointing 6: 15/10/2022 13:20-00:0023:20 (10hr) → end before the VSTP_UPDATE event
Since pointings have been agreed between the involved SOOP Coordinators, SOC will try out this change in the schedule of pointing events.
- ...
More general requests:
- EPD would like to generate an extra 800 bps (similar to LTP8 main science mode) whenever this is possible throughout LTP9.
SOC to analyse feasibility and specify the periods where extra data is available.tried out in child plan
- PHI asks whether LL flushes can be increased in volume. SOC explains that due to the recent problems with SSMM and downlink balancing, we should stay in the current concept of LL.
There are however ideas on how to get more PHI LL down during periods where this matters most:- balancing out the total load with other instrument's LL during these periods
- use of TAC in order to bring PHI data to ground quickly (note this would require Science Flushes during the TAC period, not LL flushes!)
Collisions between SOOPs and Flight/Planning events
There are cases where the switchover between SOOPs, as planned at MLP for this period, is not possible. This might be because the changeover requires a change in pointing (a SLEW) but that this inteferes with incompatible events such as a PASS. Here we list the known issues.
Telemetry situation
(TBD)